# LiD-FL: Towards List-Decodable Federated Learning

### v1

@@ 4 Analysis
In this section, we show that under proper assumptions, our algorithm converges to the optimal parameters with a large success probability. We use $X^{\prime}=\bigcup_{j\in\mathcal{H}}\{x_{i},i\in D_{j}\}$ to denote the set of all training data points possessed by honest clients. We assume that the loss function $f_{i}$ for each point $x_{i}$ is $\alpha$-strongly convex and $\beta$-smooth, the definitions of which are given in Appendix A.1. Let $w^{*}$ be the optimal parameter for the model that minimizes the global loss function $f$. To simplify the analysis, we let $\tau=1$, $\mu=0$, and $|D_{j}|=n/m$ for all $j\in[m]$. With this balanced dataset assumption, the global loss is also the empirical loss on all data points $f(w)=\frac{1}{n}\sum_{x_{i}\in X^{\prime}}f_{i}(w)$. This balanced dataset assumption can be removed by normalizing the gradient update in the algorithm.
Theorem 4.1.
Suppose $\gamma$ fraction of all clients are uncorrupted. Assume each loss function $f_{i}$ is $\alpha$-convex and $\beta$-smooth. With a list size of $q\geq\lfloor 1/\gamma\rfloor$, Algorithm 1 guarantees that after $T$ rounds, with probability at least $1-\delta T$, there is one model with parameter $w_{T}$ in the output such that
$\operatorname{\mathbf{E}}[f(w_{T})-f(w^{*})]\leq e^{-\frac{\alpha\gamma(1-\delta)}{\beta q}\cdot T}\operatorname{\mathbf{E}}[f(w_{0})-f(w^{*})]+\varepsilon.$
@@
We assume that each client independently gets an accurate estimation through such an oracle. Our analysis can be easily extended to the case where the validation processes between clients are dependent. We give an example of a validation oracle in Appendix A.2.
@@
$\leq e^{-2(p^{q+1}-1/((q+1)\gamma))k}=\delta,$
where the last inequality is due to $p^{q+1}>1/((q+1)\gamma)$ and $k\geq\ln\delta/(p^{q+1}-1/((q+1)\gamma))$.
@@
For each round $t\geq 1$, we have
$\operatorname{\mathbf{E}}[$ $(f(w_{t})-f(w^{*}))\mid F_{t}^{c}]\leq$
$\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)\operatorname{\mathbf{E}}[f(w_{t-1})-f(w^{*})]+2\eta.$
@@
Good Round: We first consider a good round. Let $w^{\prime}$ be the updated parameter sampled by the algorithm. Conditioned on the good round, the updated parameter follows a stochastic gradient descent on all data points $X^{\prime}$ possessed by all honest clients. Since $f$ is the sum of $\alpha$-convex and $\beta$-smooth functions, we have
$\operatorname{\mathbf{E}}_{w^{\prime}}[f(w^{\prime})-f(w^{*})\mid G_{t}]\leq\left(1-\frac{\alpha}{\beta}\right)\cdot(f(w_{t-1})-f(w^{*})).$
@@
$\operatorname{\mathbf{E}}_{w_{t}}[$ $f(w_{t})-f(w^{*})\mid G^{\prime}_{t}]\leq$
$\leq\left(1-\frac{\alpha}{\beta}\right)\cdot(f(w_{t-1})-f(w^{*}))+2\eta.$
@@
$\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t})-f(w^{*})\mid B^{\prime}_{t}]\leq f(w_{t-1})-f(w^{*})+2\eta.$
@@
$\leq$ $\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t})-f(w^{*})\mid G^{\prime}_{t}]\Pr\{G^{\prime}_{t}\}$
$+\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t})-f(w^{*})\mid B^{\prime}_{t}]\Pr\{B^{\prime}_{t}\}.$
@@
$\leq\left(1-\Pr\{G_{t}^{\prime}\mid F_{t}^{c}\}\cdot\frac{\alpha}{\beta}\right)\cdot(f(w_{t-1})-f(w^{*}))+2\eta$
$\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)\cdot(f(w_{t-1})-f(w^{*}))+2\eta.$
@@
$\operatorname{\mathbf{E}}[$ $(f(w_{t})-f(w^{*}))\mid F_{t}^{c}]\leq$
$\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)\operatorname{\mathbf{E}}[f(w_{t-1})-f(w^{*})]+2\eta.$
@@
$\operatorname{\mathbf{E}}[(f(w_{T})-$ $f(w^{*}))\mid F^{c}]\leq$
$\leq$ $\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)^{T}\operatorname{\mathbf{E}}[f(w_{0})-f(w^{*})]$
$+2\eta\cdot(1-\delta)\frac{\beta q}{\alpha\gamma}.$
By taking $\eta=\frac{\varepsilon\alpha\gamma}{2\beta q(1-\delta)}$, we get the conclusion. ∎
@@ 5 Experiments
In this section, we evaluate our method on image classification task with both convex and non-convex optimization objectives. Limited by space, we leave additional implementation details and full results to Appendix B.
@@ 5.1 Experimental Setup
Byzantine clients send elaborate malicious updates to the server. In LiD-FL, the server samples only one client per iteration to perform local training and obtain a model update from that client. When a Byzantine client is chosen, the honest clients do not launch local training. However, some attacks fabricate malicious updates based on the mean and variance of honest updates. In this scenario, we model an omniscient attacker (Baruch, Baruch, and Goldberg 2019), who knows the data of all clients. Thus it can mimic the statistical properties of honest updates and then craft a malicious update.
We imitate one data poison attack, the data label flipping attack (LF) and $5$ model poison attacks: the inner product manipulation attack (EPR) (Xie, Koyejo, and Gupta 2020), the random Gauss vector attack (Gauss), the Little attack (LIE) which conceals small Byzantine perturbations within the variance of honest updates (Baruch, Baruch, and Goldberg 2019), the Omniscient attack (OMN) against naive average aggregation  (Blanchard et al. 2017), and the sign flipping attack of model updates (SF). Refer to Appendix B.2 for detailed attack configurations.
Byzantine voting strategy.
@@
We compare the proposed LiD-FL algorithm with $4$ prior robust FL algorithms: naive average (FedAvg), geometric median approximated by the Weiszfeld’s algorithm (GM)  (Pillutla, Kakade, and Harchaoui 2022b) with the $1$ iteration, norm bounded mean (Norm) (Sun et al. 2019) with hyperparameter $\tau=0.215771$, and coordinate-Wise median (CWM(Yin et al. 2018).
@@
For both datasets, we simulate a distributed system with $1$ server and $35$ clients. We run experiments for $3$ levels of Byzantine fraction: $0.4$, $0.6$, and $0.8$, i.e. $14$, $21$, and $28$ out of $35$ clients are malicious.
For the FEMNIST dataset, we train a logistic regression (LR) classifier for convex optimization and a convolutional neural network (CNN) for non-convex optimization, respectively. For the CIFAR-10 dataset, we train a CNN model. Detailed model architecture and hyperparameter settings are deferred to Appendix B.3.
Once the training process is complete, each client tests all models in the global list using its local test dataset. We then calculate the average test accuracy across all clients for each model and present the highest accuracy as the test result. We repeat each experiment for $5$ times, each with a different random seed, and report the average test result across the $5$ runs  [^2] Standard deviations across runs are shown in the appendix.. The implementation is based on PRIOR [^3] https://github.com/BDeMo/pFedBreD˙public.
@@ 5.2 Analysis of Results
(a) Logistic Regression
@@
(b) Convolutional Neural Network
Table 1: Performance comparison on FEMNIST at a Byzantine fraction of 0.6.
Table 1 shows the test accuracy of various algorithms on FEMNIST with a Byzantine fraction of $0.6$. The last column displays the worst-case among the $6$ attacks. The two subtables represent the LR model and the CNN model, respectively. We use the Worst voting attack unless otherwise noted.
The results demonstrate the remarkable effectiveness of LiD-FL, as its worst test accuracy across different attacks is the highest on both LR model and CNN model while all the baseline methods collapsed. Specifically, for the LR model, the worst test accuracy of LiD-FL is $0.56$, and the other methods fail to produce a valid model. For the CNN model, none of the baselines successfully defend the learning process against the Empire, LF, LIE, OMN and SF attacks, while LiD-FL achieves a test accuracy over $72\%$ against these $5$ attacks. For the Gauss attack, the Norm shows best performance and LiD-FL achieves the second highest test accuracy. Additionally, the performance of LiD-Fl varies slightly across diverse attacks on both models, indicating the excellent robustness of LiD-FL.
@@
Figure 1 illustrates the experimental results subjected to different attack scenarios on CIFAR-10 at a Byzantine fraction of $0.6$. The results emphasize that when exposed to Empire, LF, LIE, OMN and SF attacks, LiD-FL much outperforms other methods. The test accuracy of LiD-FL is more than $0.50$ for these attacks, while the test accuracy of other algorithms all remain below $0.3$. Note that under the Gauss attack, the LiD-FL shows a lower convergence speed than baselines.
@@
Table 2 presents the test accuracy of LiD-FL under two voting attacks, where “W” and “R” represent the Worst voting attack and Random voting attack, respectively. LiD-FL achieves consistently high performance against both voting attacks, demonstrating remarkable robustness. For most attacks, the Random voting strategy is less effective than the Worst voting strategy. This is because when the new model is suboptimal but not poisoned, the Random voting scheme increases its chance of being properly optimized.
@@ 6 Conclusion
In this paper, we propose a list decodable federated learning framework called LiD-FL. Unlike traditional robust FL algorithms, LiD-FL maintains a list of global models on the server, ensuring that at least one of them is reliable in adversarial environments. One significant advantage of LiD-FL is that it has no strict restriction on the fraction of malicious clients. Our theoretical analysis proves a convergence theorem for strongly convex and smooth optimization objectives. To validate the performance of LiD-FL, we conduct numerical simulations on two image classification datasets, with both convex and non-convex loss functions. The results reveal that LiD-FL consistently outperforms previous methods under various attacks, demonstrating its effectiveness, robustness, and stability.
@@ A.1 Properties of the Loss Function
In this section, we define the $\alpha$-strongly convex and $\beta$-smooth function, to supplement the analysis in Section 4. These two properties are standard and widely used in learning algorithms (Garrigos and Gower 2024).
@@
A.2 Validation Oracle
@@ B.1 Running Environment
  • GPU: two 16GB Quadro RTX 5000 GPUs
@@ B.2 Byzantine Attacks
    Empire (EPR) (Xie, Koyejo, and Gupta 2020): the Empire attack attempts to make the inner product of the aggregated update and true update negative, which sets the corrupted updates to be the average of honest updates multiplied by $-1.1k/(m-k)$.
@@
    Gauss: the Gaussian attack replaces the model updates on corrupted clients with random Gaussian vectors with the same variance. Denote the true model update on Byzantine client $c_{j},j\in B$ by $u^{j}$, $c_{j}$ samples a vector $\tilde{u}^{j}$ from the distribution $\mathcal{N}(0,{\sigma^{j}}^{2}I)$ and sends it to the server, where $\sigma^{j}$ is the standard deviation of $u^{j}$.
@@
    Little (LIE) (Baruch, Baruch, and Goldberg 2019): the Little attack introduces tiny perturbations within the variance of honest updates. Specifically, the Byzantine clients send $u^{j}=\bar{u}-z\sigma,j\in\mathcal{B}$ to the server, where the mean and standard deviation of honest updates are $\bar{u}$, $\sigma$ respectively.
@@
    Omniscient (OMN) (Blanchard et al. 2017): the omniscient attack is designed to destroy naive average aggregation, where the mean update of all clients could be arbitrary vector $u$ by proposing $u^{j}=\frac{1}{|\mathcal{B}|}(u\cdot m-\sum_{i\in\mathcal{H}}{u^{i}}),j\in\mathcal{B}$.
@@
    Sign Flip (SF): the sign flipping attack increases the value of global loss function by sending opposite model updates. Specifically, the Byzantine client $c_{j}$ flips the sign of its true model update $u^{j}$ and submits $\tilde{u}^{j}=-u^{j}$ to the server.

### v3

@@ 4 Analysis
In this section, we show that under proper assumptions, our algorithm converges to the optimal parameters with a large success probability. We use $Z^{\prime}=\bigcup_{j\in\mathcal{H}}\{(x_{i},y_{i}),i\in D_{j}\}$ to denote the set of all training data possessed by honest clients. We assume that the loss function $f_{i}$ for each input-output pair $(x_{i},y_{i})$ is $\alpha$-strongly convex and $\beta$-smooth. The variance of the gradient noise in SGD is assumed to be bounded by $\sigma^{2}$. All assumptions are declared formally in the appendix. Let $w^{*}$ be the optimal parameter for the model that minimizes the global loss function $f$. To simplify the analysis, we let $\tau=1$, $\ell=1/\beta$ and $|D_{j}|=n/m$ for all $j\in[m]$. With this balanced dataset assumption, the global loss is also the empirical loss on $Z^{\prime}$, i.e., $f(w)=\frac{1}{|Z^{\prime}|}\sum_{(x_{i},y_{i})\in Z^{\prime}}f_{i}(w)$. This balanced dataset assumption can be removed by normalizing the gradient update in the algorithm.
@@
We assume that each client independently gets an accurate estimation through such an oracle. Our analysis can be easily extended to the case where the validation processes between clients are dependent. We give an example of a validation oracle in the appendix.
Theorem 4.2.
Suppose $\gamma$ fraction of all $m$ clients are uncorrupted. Assume each loss function $f_{i}$ is $\alpha$-stronly convex and $\beta$-smooth, and the variance of the gradient noise on each client is bounded by $\sigma^{2}$. With a list size of $q\geq\lfloor 1/\gamma\rfloor$, Algorithm 1 guarantees that after $T$ rounds, for all $m\geq\ln(1/\delta)/(2(\gamma p^{q+1}-1/((q+1))))$, there is one model with parameter $w_{T}$ in the output such that
$\operatorname{\mathbf{E}}[f(w_{T})-f(w^{*})]\leq e^{-\frac{\alpha\gamma(1-\delta)}{\beta q}\cdot T}\operatorname{\mathbf{E}}[f(w_{0})-f(w^{*})]+\varepsilon$
holds with probability at least $1-\delta T$, where $\varepsilon=q(1-\delta)(4\eta\beta+\sigma^{2})/(2\alpha\gamma)$.
@@
$\leq e^{-2(p^{q+1}-1/((q+1)\gamma))k}\leq\delta,$
where the last inequality is due to $p^{q+1}>1/((q+1)\gamma)$ and $k\geq\ln(1/\delta)/(2(p^{q+1}-1/((q+1)\gamma)))$.
@@
Let $\check{\eta}=2\eta+\sigma^{2}/(2\beta)$, for each round $t\geq 1$, we have
$\operatorname{\mathbf{E}}[(f(w_{t})-f(w^{*}))\mid F_{t}^{c}]\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)\operatorname{\mathbf{E}}[f(w_{t-1})-f(w^{*})]+\check{\eta},$
@@
Good Round: We first consider a good round. Let $w^{\prime}$ be the updated parameter sampled by the algorithm. Conditioned on the good round, the updated parameter follows a stochastic gradient descent on all data possessed by honest clients $Z^{\prime}$. Since $f$ is the sum of $\alpha$-convex and $\beta$-smooth functions, we have
$\operatorname{\mathbf{E}}_{w^{\prime}}[f(w^{\prime})-f(w^{*})\mid G_{t}]\leq\left(1-\frac{\alpha}{\beta}\right)\cdot\operatorname{\mathbf{E}}_{w^{\prime}}[(f(w_{t-1})-f(w^{*}))]+\sigma^{2}/(2\beta).$
@@
$\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t})-f(w^{*})\mid G^{\prime}_{t}]\leq\left(1-\frac{\alpha}{\beta}\right)\cdot\operatorname{\mathbf{E}}_{w_{t}}[(f(w_{t-1})-f(w^{*}))]+\check{\eta}.$
@@
$\operatorname{\mathbf{E}}_{w_{t}}[$ $f(w_{t})-f(w^{*})\mid B^{\prime}_{t}]\leq\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t-1})-f(w^{*})]+2\eta.$
@@
$\leq$ $\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t})-f(w^{*})\mid G^{\prime}_{t}]\Pr\{G^{\prime}_{t}\}+\operatorname{\mathbf{E}}_{w_{t}}[f(w_{t})-f(w^{*})\mid B^{\prime}_{t}]\Pr\{B^{\prime}_{t}\}.$
@@
$\leq\left(1-\Pr\{G_{t}^{\prime}\mid F_{t}^{c}\}\cdot\frac{\alpha}{\beta}\right)\operatorname{\mathbf{E}}_{w_{t}}[(f(w_{t-1})-f(w^{*}))]+\check{\eta}$
$\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)\operatorname{\mathbf{E}}_{w_{t}}[(f(w_{t-1})-f(w^{*}))]+\check{\eta}.$
@@
$\operatorname{\mathbf{E}}[(f(w_{t})-f(w^{*}))\mid F_{t}^{c}]\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)\operatorname{\mathbf{E}}[f(w_{t-1})-f(w^{*})]+\check{\eta}.$
@@
$\operatorname{\mathbf{E}}[(f(w_{T})-f(w^{*}))\mid F^{c}]\leq\left(1-\frac{\alpha\gamma}{\beta q}(1-\delta)\right)^{T}\operatorname{\mathbf{E}}[f(w_{0})-f(w^{*})]+\check{\eta}\cdot(1-\delta)\frac{\beta q}{\alpha\gamma}.$
By taking $\check{\eta}=\frac{\varepsilon\alpha\gamma}{\beta q(1-\delta)}$, we get the conclusion. ∎
@@ 5 Experiments
In this section, we evaluate our method on image classification tasks with both convex and non-convex optimization objectives. Limited by space, we leave additional implementation details and full results to the appendix.
@@ 5.1 Experimental Setup
Byzantine clients send elaborate malicious updates to the server. In LiD-FL, the server samples only one client per iteration to perform local training and obtain a model update from that client. When a Byzantine client is chosen, the honest clients do not launch local training. However, some attacks fabricate malicious updates based on the mean and variance of honest updates. In this scenario, we model an omniscient attacker (Baruch et al. 2019), who knows the data of all clients. Thus it can mimic the statistical properties of honest updates and then craft a malicious update.
We simulate one data poison attack, the data label flipping attack (LF) and $5$ model poison attacks: the inner product manipulation attack (EPR) (Xie et al. 2020), the random Gauss vector attack (Gauss), the little perturbations attack (LIE) (Baruch et al. 2019), the Omniscient attack (OMN)  (Blanchard et al. 2017), and the sign flipping attack (SF). Refer to the appendix for detailed attack configurations.
@@
We compare the proposed LiD-FL algorithm with $4$ prior robust FL algorithms: naive average (FedAvg(McMahan et al. 2017), geometric median approximated by the Weiszfeld’s algorithm (GM)  (Pillutla et al. 2022b) with the $1$ iteration, norm bounded mean (Norm) (Sun et al. 2019) with hyperparameter $\tau=0.215771$, and coordinate-wise median (CWM(Yin et al. 2018). In experiments we incorporate the notion of momentum in local updates. We denote the initial momentum and momentum coefficient of client $c_{j}$ by $v_{0}^{(j)}$ and $\mu$, respectively. The local training step with momentum is given by $v_{r+1}^{(j)}\leftarrow\mu\cdot v_{r}^{(j)}+g_{r}^{(j)}$, followed by $w_{r+1}^{(j)}\leftarrow w_{r}^{(j)}-\ell\cdot v_{r+1}^{(j)}$.
@@
For both datasets, we simulate a distributed system with $1$ server and $35$ clients. We perform experiments for $3$ levels of Byzantine fraction: $0.4$, $0.6$, and $0.8$. For the FEMNIST dataset, we train a logistic regression (LR) classifier for convex optimization and a convolutional neural network (CNN) for non-convex optimization, respectively. For the CIFAR-10 dataset, we train a CNN model. Detailed model architecture and hyperparameter settings are deferred to the appendix.
Once the training process is complete, each client tests all models in the global list using its local test dataset. We then calculate the average test accuracy across all clients for each model and present the highest accuracy as the test result. We repeat each experiment for $5$ times, each with a different random seed, and report the average test result across the $5$ runs  [^3] Standard deviations across runs are shown in the appendix.. The implementation is based on PRIOR [^4] https://github.com/BDeMo/pFedBreD_public.
@@ 5.2 Analysis of Results
Table 1: Performance comparison on FEMNIST at a Byzantine fraction of 0.6, with the LR global model.
@@
Table 2: Performance comparison on FEMNIST at a Byzantine fraction of 0.6, with the CNN global model.
Table 1 and Table 2 show the test accuracy of various algorithms on FEMNIST at a Byzantine fraction of $0.6$ with the LR model and the CNN model, respectively. The last column displays the worst-case among the $6$ attacks. We use the Worst voting attack unless otherwise noted.
The results demonstrate the remarkable effectiveness of LiD-FL, as its worst test accuracy across different attacks is the highest on both LR model and CNN model while all the baseline methods collapsed. Specifically, for the LR model, the worst test accuracy of LiD-FL is $0.56$, and the other methods fail to produce a valid model. For the CNN model, none of the baselines successfully defend the learning process against the EPR, LF, LIE, OMN and SF attacks, while LiD-FL achieves a test accuracy over $72\%$ against these $5$ attacks. For the Gauss attack, the Norm shows best performance and LiD-FL achieves the second highest test accuracy. Additionally, the performance of LiD-FL varies slightly across diverse attacks on both models, indicating the excellent robustness of LiD-FL.
@@
Figure 1 illustrates the experimental results on CIFAR-10 at a Byzantine fraction of $0.6$. The results emphasize that when exposed to EPR, LF, LIE, OMN and SF attacks, LiD-FL much outperforms other methods. The test accuracy of LiD-FL is more than $0.50$ for these attacks, while the test accuracy of other algorithms all remain below $0.3$. Note that under the Gauss attack, the LiD-FL shows a lower convergence speed than baselines.
@@
Table 3 presents the test accuracy of LiD-FL under two voting attacks, where “W” and “R” represent the Worst voting attack and Random voting attack, respectively. LiD-FL achieves consistently high performance against both voting attacks, demonstrating remarkable robustness. In most cases, the Random voting attack is less effective than the Worst voting attack. This is because when the new model is suboptimal but not poisoned, the Random voting scheme increases its chance of being optimized.
@@ 6 Conclusion
In this paper, we propose a list decodable federated learning framework called LiD-FL. Unlike traditional robust FL algorithms, LiD-FL maintains a list of global models on the server, ensuring that at least one of them is reliable in adversarial environments. One significant advantage of LiD-FL is that it has no strict restriction on the fraction of malicious clients. Our theoretical analysis proves a convergence theorem for strongly convex and smooth optimization objectives. Numerical simulations on two datasets demonstrate the effectiveness, robustness, and stability of LiD-FL.
@@ A.1 Properties of the Loss Function
In this section, we define the $\alpha$-strongly convex and $\beta$-smooth function, to supplement the analysis in Section 4 of the main paper. These two properties are standard and widely used in learning algorithms (Garrigos and Gower 2024).
@@
A.2 Properties of the Gradient Noise
We assume that the gradient noise in SGD has bounded variance to support our analysis, an assumption widely adopted in machine learning research (Ghadimi and Lan 2013; Bottou et al. 2018; Farhadkhani et al. 2022).
Definition A.3.
Bounded variance of gradient noise. For any parameter $\theta$, the gradient noise at $\theta$ in SGD process is $\xi(\theta)$. There exists $\sigma\geq 0$ satisfying
$\operatorname{\mathbf{E}}[\xi(\theta)]=0,\mathrm{Var}(\xi(\theta))\leq\sigma^{2}.$
A.3 Validation Oracle
@@ B.1 Running Environment
  • GPU: two 16GB Quadro RTX 5000 GPUs and four NVIDIA GeForce RTX 4090 GPUs
@@ B.2 Byzantine Attacks
    Empire (EPR) (Xie et al. 2020): the Empire attack attempts to make the inner product of the aggregated update and true update negative, which sets the corrupted updates to be the average of honest updates multiplied by $-1.1k/(m-k)$.
@@
    Gauss: the Gaussian attack replaces the model updates on corrupted clients with random Gaussian vectors with the same variance. Denote the true model update on Byzantine client $c_{j},j\in B$ by $u^{(j)}$, $c_{j}$ samples a vector $\tilde{u}^{(j)}$ from the distribution $\mathcal{N}(0,{\sigma^{(j)}}^{2}I)$ and sends it to the server, where $\sigma^{(j)}$ is the standard deviation of $u^{(j)}$.
@@
    Little (LIE) (Baruch et al. 2019): the Little attack introduces tiny perturbations within the variance of honest updates. Specifically, the Byzantine clients send $\tilde{u}^{(j)}=\bar{u}-z\sigma,j\in\mathcal{B}$ to the server, where the mean and standard deviation of honest updates are $\bar{u}$, $\sigma$ respectively.
@@
    Omniscient (OMN) (Blanchard et al. 2017): the omniscient attack is designed to destroy naive average aggregation, where the mean update of all clients could be arbitrary vector $u$ by proposing $\tilde{u}^{(j)}=\frac{1}{|\mathcal{B}|}(u\cdot m-\sum_{i\in\mathcal{H}}{u^{(i)}}),j\in\mathcal{B}$.
@@
    Sign Flip (SF): the sign flipping attack increases the value of global loss function by sending opposite model updates. Specifically, the Byzantine client $c_{j}$ flips the sign of its true model update $u^{(j)}$ and submits $\tilde{u}^{(j)}=-u^{(j)}$ to the server.